OpenStudy (daggiemib):

A spotlight is mounted on a wall 7.4 feet above the floor in an office building. It is used to light a door 9.3 feet from the wall. To the nearest degree, what is the angle of depression from the spotlight to the bottom of the door? 39 degrees 51 degrees 53 degrees 37 degrees

OpenStudy (anonymous):

I belive you use SOH-CAH-TOA

OpenStudy (daggiemib):

toa i got 51 but i wanted to check if its correct

OpenStudy (anonymous):

you are right

OpenStudy (daggiemib):

yay thanks:)))

OpenStudy (anonymous):

That's what I got :)

OpenStudy (anonymous):

me too

OpenStudy (daggiemib):

the answer was 39 degrees.... i dont know how but i got it wrong:(

OpenStudy (daggiemib):

@imqwerty can you explain how i should have done this?

imqwerty (imqwerty):

:)

imqwerty (imqwerty):

 so theta is our angle of depression we can say that\[\angle(CAB)=\theta \]because theta and angle(CAB) are corresponding angles now we need to find this theta \[\tan(CAB)=\frac{ BC }{ AB }\]\[\tan(\theta)=\frac{ 7.4 }{ 9.3 }\]\[\tan(\theta)=0.7956\]\[\theta=\tan^{-1} (0.7956)\]\[\theta=39^{o}\]